emolyrics2:

f(x) = x2 – 4x – 21.

carmelle:

What is the question being asked?

carmelle:

@emolyrics2 - you can just put the question here so someone can help :)

axie:

I think its solve for f(x)

steevh:

f(x) = -2x-21 = 0 f(x) = -2x = 21 f(x) = x = 21/2 f(x) = 0 when x = 10.5

brianagatica14:

@steevh wrote:
f(x) = -2x-21 = 0 f(x) = -2x = 21 f(x) = x = 21/2 f(x) = 0 when x = 10.5
god jog? ig

brianagatica14:

@brianagatica14 wrote:
@steevh wrote:
f(x) = -2x-21 = 0 f(x) = -2x = 21 f(x) = x = 21/2 f(x) = 0 when x = 10.5
god jog? ig
it's kinda confusing but ok

steevh:

you can switch the value f(x) with 0 so you start with x2 – 4x – 21 = 0 and so on